OpenStudy (anonymous):

simplify 10*3sqrt 7 - 7*4sqrt 7 - 9*3sqrt 7 + 4sqrt 7

OpenStudy (anonymous):

0

OpenStudy (anonymous):

\[\LARGE 10 \sqrt[3]{7} - 7\sqrt[4]{7} - 9\sqrt[3]{7} + \sqrt[4]{7}\]

OpenStudy (anonymous):

\[10x-7y-9x+y=x-6y\]

OpenStudy (anonymous):

i.e. combine like terms, you get \[\sqrt[3]{7}-6\sqrt[4]{7}\]